A child starts sledding from rest, going down a 40-m-long 7.5° incline, then coasting across a horizontal stretch. The mass of sled + child is 35 kg, and the coefficient of kinetic friction is 0.060.
Determine how much energy that is transferred to thermal energy.
1
Expert's answer
2020-12-14T12:15:39-0500

"E = E_p+E_k+Q"

"\text{where:}"

"E \text { energy, }E_p \text { potential energy,}"

"E_k \text { kinetic energy, }Q \text{ internal energy}"

"\text{at the top of the slope:}"

"E_1 = E_{p1}+ E_{k1}+Q_1= E_{p1}+Q_1"

"E_{k1}= 0 \text{ as }V= 0\text{ (body speed)}"

"\text{the sled has finished moving:}"

"E_2 = E_{p2}+ E_{k2}+Q_2= E_{p2}+Q_2"

"E_{k2}= 0 \text{ as }V= 0\text{ (body speed)}"

"\text{No action of outside forces according to the law of conservation of energy:}"

"E_1=E_2;Q_2-Q_1=E_{p1}-E_{p2}= mg\Delta{}h"

"\Delta{h}= l *sin\alpha=40*\sin{7.5^0}= 5.22\text{ slope height}"

"mg\\Delta{}h=35*9.8*5,22=1\u00a0790.46 J"

Answer: 1 790.46 J energy that is transferred to thermal energy